OpenStudy (anonymous):

evaluate the expression below when x = 3. 5x2 - 12x + 3

OpenStudy (anonymous):

The possible answers are 1.) 3 2.) 11/3 3.) 11/2 4.) 26/3

OpenStudy (igreen):

\(5x^2 - 12x + 3\) Input 3 for x: \(5(3)^2 - 12(3) + 3\) The exponent is over the 3, so we have to simplify \(3^2\). \(5(9) - 12(3) + 3\) 5 * 9 = 45 \(45 - 12(3) + 3\) -12 * 3 = -36 \(45 - 36 + 3\) 45 - 36 = 9 \(9 + 3\) 9 + 3 = 12 \(12\)

OpenStudy (igreen):

I did something wrong.

OpenStudy (igreen):

But it seems correct..

OpenStudy (igreen):

That's correct.. the answer is 12.. right, @mathmale ?

OpenStudy (igreen):

@JulianoRossiel

OpenStudy (anonymous):

No the following are 1.) 3 2.) 11/3 3.) 11/2 4.) 26/3

OpenStudy (igreen):

Are you sure this is the expression? @JulianoRossiel \(5x^2 - 12x + 3\)

OpenStudy (mathmale):

I like the approach 'iGreen' has taken here: He/she has replaced every occurrence of 'x' with parentheses enclosing the x value: x => (3). Note how your 'x2' (which should be written 'x^2' becomes (3)^2, or 9.

OpenStudy (mathmale):

'x squared' can be written as x^2, \[x^2\] or

OpenStudy (igreen):

I have made no mathematical errors..

OpenStudy (mathmale):

but preferably not 'x2'.

OpenStudy (igreen):

Yet my answer is not one of the options.

OpenStudy (mathmale):

evaluate the expression below when x = 3. 5x2 - 12x + 3: or, evaluate 5x^2 - 12x + 3 for x = 3: 5(3)^2 - 12(3) + 3 = 5(9) -36 + 3 = ???

OpenStudy (igreen):

5(9) - 36 + 3 5 * 9 = 45 45 - 36 + 3 9 + 3 12

OpenStudy (igreen):

However, 12 is not an option. 12 ≠ 3 12 ≠ 11/3 12 ≠ 11/2 12 ≠ 26/3

OpenStudy (anonymous):

5x^2-12/x+3 Sorry I wrote the equation wrong

OpenStudy (mathmale):

Uh-oh. Please replace each occurrence of 'x' with '(x)' now. Then evaluate the expression.

OpenStudy (anonymous):

I need help the expression

OpenStudy (igreen):

\(\dfrac{5x^2-12}{x+3}\) Input 3 in for every \(x\) term: \(\dfrac{5(3)^2-12}{(3)+3}\) Simplify the exponent: \(\dfrac{5(9)-12}{3+3}\) Multiply: \(\dfrac{45-12}{3+3}\) Subtract: \(\dfrac{33}{3+3}\) Add: \(\dfrac{33}{6}\) Simplify the fraction: \(\dfrac{11}{2}\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @JulianoRossiel

OpenStudy (anonymous):

I will and thanks for your help